9. Recall that Babe’s Bats faces an inverse demand curve of P = 25-2.5Q and marginal cost curve MC = 5Q. Calculate the deadweight loss from market power at the firm’s profit-maximizing level of output.
Answer #1

Answer

the firm produces at MR=MC

MR=25-5Q .......... An MR curve is double sloped than an inverse linear demand

MC=5Q

equating both equations

5Q=25-5Q

10Q=25

Q=2.5

P=25-2.5*2.5=18.75

MC=5*2.5=12.5

==============

A social efficient output is at MC=P

5Q=25-2.5Q

7.5Q=25

Q=3.33333333

===========

deadweight loss =0.5* change in quantity *(P-MC)

=0.5*(3.333333-2.5)*(18.75-12.5)

=2.60416563

=2.60

the DWL is $2.60